# Imaginarity of quantum channels: Refinement and Alternative

### v1

@@ 4 Imaginarity measures of quantum channels
The Choi theorem [27] tell us that the Choi matrix of the channel $\phi$ is positive semidefinite, so that $\chi_{\phi}\geqslant 0$, we can see that, for any quantum channel $\phi$ with $\phi(\cdot)=\sum_{n}K_{n}\cdot K_{n}^{\dagger}$, its Choi state $\chi_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|$, thus the trace:
$\text{Tr}(\chi_{\phi})=\text{Tr}\Big{[}\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|\Big{]}=\frac{1}{|A|}\sum_{j}\text{Tr}\Big{[}\phi(|j\rangle\langle j|)\Big{]}=\frac{1}{|A|}\sum_{j}1=1.$
In this way, we can embed Choi states into quantum states:
$\mathcal{C}_{AB}\hookrightarrow\mathcal{D}_{AB}.$
It seems then that many imaginarity measures for quantum states can be equally applied to quantum channels. Thus in this section, we introduce three imaginarity measures of quantum channels, that is, measures via on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively. We can exactly verify these three measures in two ways.
4.1 Measure via on the robustness
The Measure via on the robustness is appeared in many resource theories [1, 18]. Similar to these, we define the imaginarity measure of quantum channels as follow.
@@ 4.1 Measure via on the robustness
$C_{r}(\phi)$ is a proper imaginarity measure of quantum channels.
@@
We only verify $\bm{(C_{1})}$+$\bm{(C_{3})}$+$\bm{(C_{4})}$ since $\bm{(C_{2})}$ can be derived.
@@
Obviously $C_{r}(\phi)\geqslant 0$. If $C_{r}(\phi)=0$, that means $\frac{\chi_{\phi}+0}{1+0}=\chi_{\psi}$, that is $\phi=\psi\in\mathcal{RC}_{AB}$. The converse is certainly true.
@@
Let $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, for any quantum channel $\phi$ with its pseudomixture, under the requirement of $\bm{(C_{3})}$, $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$ and $\chi_{\phi_{m}}=M_{m}\chi_{\phi}M_{m}^{\dagger}/p_{m}$. Then we obtain
$M_{m}\chi_{\phi}M_{m}^{\dagger}=[1+C_{r}(\chi_{\phi})]M_{m}\chi_{\phi_{1}}M_{m}^{\dagger}-C_{r}(\chi_{\phi})M_{m}\chi_{\phi_{2}}M_{m}^{\dagger}.$
@@
We get $\chi_{\phi_{m}}=(1+s_{m})\chi_{\phi_{1}}^{m}-s_{m}\chi_{\phi_{2}}^{m}$, that is, $\frac{\chi_{\phi_{m}}+s_{m}\chi_{\phi_{2}}^{m}}{1+s_{m}}=\chi_{\phi_{1}}^{m}$. Thus we get $C_{r}(\chi_{\phi_{m}})\leqslant s_{m}$ from definition, that is, $p_{m}C_{r}(\chi_{\phi_{m}})\leqslant C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$.
Therefore, we obtain
$\sum_{m}p_{m}C_{r}(\chi_{\phi_{m}})$$\leqslant$ $\sum_{m}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$
@@
For simplicity, we only consider two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\phi=p\phi_{1}+(1-p)\phi_{2}$ with any $p\in(0,1)$, that is, $\chi_{\phi}=p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}}$. The measurements $C_{r}$ for $\phi_{1}$ and $\phi_{2}$ are $C_{r}(\chi_{\phi_{1}})$ and $C_{r}(\chi_{\phi_{2}})$. Let $s_{p}=pC_{r}(\chi_{\phi_{1}})+(1-p)C_{r}(\chi_{\phi_{2}})$.
@@
We have $\chi_{\phi}=(1+s_{p})\chi_{\psi}-s_{p}\chi_{\tilde{\phi}}$, that is, $\chi_{\psi}=\frac{\chi_{\phi}+s_{p}\chi_{\tilde{\phi}}}{1+s_{p}}$, thus we get $s_{p}\geqslant C_{r}(\chi_{\phi})$.
@@
4.2 Measure via on the trace norm
To obtain the channel is real, we can consider the trace norm:
@@ 4.2 Measure via on the trace norm
where $\chi_{\phi}^{*}$ means the conjugate of $\chi_{\phi}$ and $|\chi_{\phi}-\chi_{\phi}^{*}|=\sqrt{(\chi_{\phi}-\chi_{\phi}^{*})^{\dagger}(\chi_{\phi}-\chi_{\phi}^{*})}$.
@@
There is $||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}\geqslant||\Theta(\chi_{\phi})-\Theta(\chi_{\phi}^{*})||_{\text{Tr}}$ for any CPTP map [31], and it certainly holds for real superchannels. Since $\Theta$ is a real superchannel, we can know that $[\Theta(\chi_{\phi})]^{*}=\Theta(\chi_{\phi}^{*})$ due to
$[\Theta(\chi_{\phi})]^{*}=\left[\sum_{m}M_{m}\chi_{\phi}M_{n}^{\dagger}\right]^{*}=\sum_{m}M_{m}\chi_{\phi}^{*}M_{n}^{\dagger}=\Theta(\chi_{\phi}^{*}).$
@@
$C_{t}(\chi_{\phi})$ $=$ $C_{t}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}||_{\text{Tr}}$
$=$ $||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}^{*}\oplus p_{2}\chi_{\phi_{2}}^{*})||_{\text{Tr}}=||p_{1}(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})\oplus p_{2}(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
$=$ $p_{1}||(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})||_{\text{Tr}}+p_{2}||(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}=p_{1}C_{t}(\chi_{\phi_{1}})+p_{2}C_{t}(\chi_{\phi_{2}}).$
@@
4.3 Measure via on the sandwiched Rényi relative entropy
The sandwiched Rényi relative entropy [32] defined for quantum states with the parameter $\alpha$ as follow:
@@ 4.3 Measure via on the sandwiched Rényi relative entropy
where Now, we will introduce the imaginarity measure of quantum channels via on the sandwiched Rényi relative entropy, that is,
@@
where $\frac{1}{2}\leqslant\alpha<1$. We have the following results.
@@
It is shown [32] that $D_{\alpha}(\rho||\sigma)\geqslant 0$ when $\frac{1}{2}\leqslant\alpha<1$ and equation holds if and only if $\rho=\sigma$, that is, $\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha}\leqslant 1$. Thus we consider Choi state as a quantum state and then we have, $C_{\alpha}(\phi)\geqslant 0$. The equation holds if and only if $\chi_{\phi}=\chi_{\phi}^{*}$. The condition $\bm{(C_{1})}$ is satisfied.
@@
$\iff$ log $\text{log }\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{log }\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{Tr}\left\{([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
@@
$\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$ $=$ $\text{Tr}\left\{\left[(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}\right]^{\frac{1-\alpha}{2\alpha}}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\left[(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}\right]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
@@
$C_{\alpha}(\phi)=1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}=\sum_{j=1}^{2}p_{j}\left\{1-\text{Tr}\left[(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\right\}=p_{1}C_{\alpha}(\phi_{1})+p_{2}C_{\alpha}(\phi_{2})$
@@ 5 Conclusion
We refine the framework of imaginarity of quantum channels which originally only has $\bm{(C_{1})}$ and $\bm{(C_{2})}$ for its measure. We add two conditions, strong monotonicity $\bm{(C_{3})}$ and convexity $\bm{(C_{4})}$. Besides, we give some properties of imaginarity measure of quantum channels. Sometimes it is not easy to verify strong monotonicity and convexity, so we introduce an alternative framework, that is, $\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when $\bm{(C_{1})}$ and $\bm{(C_{2})}$ hold. This is analogous to imaginarity or coherence of quantum states. Since Choi states can embed quantum states, we also think that imaginarity measures of quantum states can also be used for quantum channels and present three imaginarity measures for quantum channels, they are $C_{r}$, $C_{t}$ and $C_{\alpha}$ via on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively.

### v2

@@ 4 Imaginarity measures of quantum channels
In this section, we introduce three imaginarity measures of quantum channels, that is, measures based on the robustness, the trace norm, and the sandwiched Rényi relative entropy, respectively.
4.1 Measure based on the robustness
The Measure based on the robustness has appeared in many resource theories [1, 18]. Similarly, we define the imaginarity measure of quantum channels as follows.
@@ 4.1 Measure based on the robustness
$C_{r}(\phi)$ is an imaginarity measure of quantum channels.
@@
We only verify $\bm{(C_{1})}$+$\bm{(C_{3})}$+$\bm{(C_{4})}$ since $\bm{(C_{2})}$ can be derived by $\bm{(C_{3})}$ and$\bm{(C_{4})}$.
@@
It can be observed that $C_{r}(\phi)\geqslant 0$. When $C_{r}(\phi)=0$, this implies $\frac{\chi_{\phi}+0}{1+0}=\chi_{\psi}$, which means $\phi=\psi\in\mathcal{RC}_{AB}$. The converse is also clearly true.
@@
Let $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$. For any quantum channel $\phi$ with its pseudomixture, under the requirement of $\bm{(C_{3})}$, we have $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$ and $\chi_{\phi_{m}}=M_{m}\chi_{\phi}M_{m}^{\dagger}/p_{m}$. Then, we obtain:
$M_{m}\chi_{\phi}M_{m}^{\dagger}$
$=$ $[1+C_{r}(\chi_{\phi})]M_{m}\chi_{\phi_{1}}M_{m}^{\dagger}-C_{r}(\chi_{\phi})M_{m}\chi_{\phi_{2}}M_{m}^{\dagger}.$
@@
We obtain $\chi_{\phi_{m}}=(1+s_{m})\chi_{\phi_{1}}^{m}-s_{m}\chi_{\phi_{2}}^{m}$. This means $\frac{\chi_{\phi_{m}}+s_{m}\chi_{\phi_{2}}^{m}}{1+s_{m}}=\chi_{\phi_{1}}^{m}$. Thus we get $C_{r}(\chi_{\phi_{m}})\leqslant s_{m}$ from definition, that is, $p_{m}C_{r}(\chi_{\phi_{m}})\leqslant C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$.
Therefore, we have:
$\sum_{m}p_{m}C_{r}(\chi_{\phi_{m}})$
$\leqslant$ $\sum_{m}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$
@@
For simplicity, we consider only two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\phi=p\phi_{1}+(1-p)\phi_{2}$ with any $p\in(0,1)$. This leads to $\chi_{\phi}=p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}}$. Let $s_{p}=pC_{r}(\chi_{\phi_{1}})+(1-p)C_{r}(\chi_{\phi_{2}})$.
@@
We have $\chi_{\phi}=(1+s_{p})\chi_{\psi}-s_{p}\chi_{\tilde{\phi}}$, that is, $\chi_{\psi}=\frac{\chi_{\phi}+s_{p}\chi_{\tilde{\phi}}}{1+s_{p}}$, thus we obtain $s_{p}\geqslant C_{r}(\chi_{\phi})$.
@@
4.2 Measure based on the trace norm
The trace norm,
@@ 4.2 Measure based on the trace norm
can be used to construct a measure, where $\chi_{\phi}^{*}$ means the conjugate of $\chi_{\phi}$ and $|\chi_{\phi}-\chi_{\phi}^{*}|=\sqrt{(\chi_{\phi}-\chi_{\phi}^{*})^{\dagger}(\chi_{\phi}-\chi_{\phi}^{*})}$.
@@
Due to $||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}\geqslant||\Theta(\chi_{\phi})-\Theta(\chi_{\phi}^{*})||_{\text{Tr}}$ for any CPTP map [30], it certainly holds for real superchannels. Since $\Theta$ is a real superchannel, it can be seen that $[\Theta(\chi_{\phi})]^{*}=\Theta(\chi_{\phi}^{*})$ due to
$[\Theta(\chi_{\phi})]^{*}$
$=$ $\left[\sum_{m}M_{m}\chi_{\phi}M_{n}^{\dagger}\right]^{*}$
$=$ $\sum_{m}M_{m}\chi_{\phi}^{*}M_{n}^{\dagger}$
$=$ $\Theta(\chi_{\phi}^{*}).$
@@
$C_{t}(\chi_{\phi})$
$=$ $C_{t}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})$
$=$ $||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}||_{\text{Tr}}$
$=$ $||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}^{*}\oplus p_{2}\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
$=$ $||p_{1}(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})\oplus p_{2}(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
$=$ $p_{1}||(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})||_{\text{Tr}}+p_{2}||(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
$=$ $p_{1}C_{t}(\chi_{\phi_{1}})+p_{2}C_{t}(\chi_{\phi_{2}}).$
@@
4.3 Measure based the sandwiched Rényi relative entropy
The sandwiched Rényi relative entropy [31] is defined for quantum states with the parameter $\alpha$ as follows:
@@ 4.3 Measure based the sandwiched Rényi relative entropy
Now, we will introduce the imaginarity measure of quantum channels based on the sandwiched Rényi relative entropy, defined as:
@@
where $\frac{1}{2}\leqslant\alpha<1$. We have the following results:
@@
It is shown [31] that $D_{\alpha}(\rho||\sigma)\geqslant 0$ when $\frac{1}{2}\leqslant\alpha<1$, and equation holds if and only if $\rho=\sigma$, which implies that $\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha}\leqslant 1$. Thus, considering Choi state as a quantum state, we find that $C_{\alpha}(\phi)\geqslant 0$. The equation holds if and only if $\chi_{\phi}=\chi_{\phi}^{*}$, satisfying the condition $\bm{(C_{1})}$.
@@
$\iff$ log $\text{log }\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
$\leqslant\text{log }\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
$\leqslant\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
$\leqslant\text{Tr}\left\{([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
@@
$\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
@@
$C_{\alpha}(\phi)$
$=$ $1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
$=$ $\sum_{j=1}^{2}p_{j}\left\{1-\text{Tr}\left[(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\right\}$
$=$ $p_{1}C_{\alpha}(\phi_{1})+p_{2}C_{\alpha}(\phi_{2})$
@@ 5 Conclusion
We introduce the framework of imaginarity of quantum channels, and introduce an alternative framework. Besides, we give some properties of imaginarity measure of quantum channels. We also give three imaginarity measures for quantum channels. They are $C_{r}$, $C_{t}$ and $C_{\alpha}$ based on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively.